OpenStudy (anonymous):

find the coordinates for the foci. x*2/16 + y*2/81 = 1

OpenStudy (anonymous):

Formula for the focal distance of an ellipse a^2 = b^2 + c^2 81 = 16 + c^2 c^2 = 65 c = √65 Since this is a vertical oriented ellipse, have the focii on the major axis. (0, √65), (0, -√65)